A 1.0-cm-diameter pipe widens to 2.0 cm, then narrows to 0.5 cm. Liquid flows through the first segment at a speed of 9.0 m/s a. what is the speed in the second segment?

b. What is the speed in the third segment?

c. what is the volume flow rate through the pipe?
Answer #1

Solution)

Part a)

A1V1 = A2 V2

π.01^2/4 *9 = π .02^2/4 * v

v = 2.25 m/s

b) A3V3 = A2 V2

π.01^2/4 *9 = π .005^2/4 * v

v = 36 m/s

c)

Q = A1V1 = π.01^2/4 *9 = 0.0007065 L/s (Ans)

======

Good luck!:)